1 0 -2 -6 ? Solve this
Answer by rothauserc(4718)

1, 0, -2, -6
:
a(n) = 2 - 2^n
:
a(0) = 2 - 2^0 = 1
a(1) = 2 - 2^1 = 0
a(2) = 2 - 2^2 = -2
a(3) = 2 - 2^3 = -6
:
a(4) = 2 - 2^4 = -14
:
*****************
1, 0, -2, -6, -14
*****************
: